OpenStudy (yuii):

After 8 years, what is the total amount of a compound interest investment of $25,000 at 3% interest, compounded quarterly? $25,377.11 $25,759.92 $26,148.50 $31,752.78

OpenStudy (yuii):

I think it's D.

OpenStudy (isaiahhlindsayy):

Best Answer: PV = $25,000 n = number of periods i = interest rate per period FV = Future Value Compound interest formula: FV = PV(1+i)^n FV = $25,000(1+.03/4)^32 FV = $31,752.78 Eggolas M · 5 years

OpenStudy (isaiahhlindsayy):

d

OpenStudy (yuii):

Thank you.